# Sensipy: simulate gamma-ray observations of transient astrophysical sources

### v1

@@ 5 Cherenkov Telescope Array Observatory, Bologna, Italy
6 Institut de Fisica d’Altes Energies (IFAE), Bellaterra (Barcelona), Spain
@@
  • How can intrinsic source properties (eg distance, flux) and observing conditions (eg latency, telescope pointing) affect detectability?
@@
Given these inputs, sensipy builds upon primitives provided by astropy and gammapy to provide the following main functionalities (The Astropy and Price-Whelan Collaboration et al. 2022; Donath et al. 2023). In addition, mock datasets are provided with working code examples, and batteries are included for easy access to publicly-available IRFs, eg (Cherenkov Telescope Array Observatory and Consortium 2021).
@@
Such workflows based on sensipy modules are already being internally evaluated within the MAGIC, Large-Size Telescope (LST), and CTAO collaborations for follow-up of both GW and GRB alerts (eg, Green et al. 2024; Patricelli et al. 2022).
@@
AI tools (GitHub Copilot, Grammarly) were used to proofread documentation and pull requests. These AI-assisted suggestions were carefully reviewed and approved by the authors of this manuscript. AI tools were not used in the writing of this manuscript in any capacity.

### v2

@@ 5 Cherenkov Telescope Array Observatory, Bologna, Italy
6 Institut de Física d’Altes Energies (IFAE), Bellaterra (Barcelona), Spain
@@
  • How can intrinsic source properties (e.g. distance, flux) and observing conditions (e.g. latency, telescope pointing) affect detectability?
@@
Given these inputs, sensipy builds upon primitives provided by astropy and gammapy to provide the following main functionalities (The Astropy and Price-Whelan Collaboration et al. 2022; Donath et al. 2023). In addition, mock datasets are provided with working code examples, and batteries are included for easy access to publicly-available IRFs, e.g. (Cherenkov Telescope Array Observatory and Consortium 2021).
@@
Such workflows based on sensipy modules are already being internally evaluated within the MAGIC, Large-Size Telescope (LST), and CTAO collaborations for follow-up of both GW and GRB alerts (e.g., Green et al. 2024; Patricelli et al. 2022).
@@
AI tools (GitHub Copilot, Grammarly) were used to proofread documentation and pull requests as well as scaffold tests. Instances where functions or classes were primarily generated with assistance from GitHub Copilot (auto model) are explicitly marked in the Python docstrings. All AI-assisted suggestions were carefully reviewed and approved by the authors of this manuscript. AI tools were not used in the writing of this manuscript in any capacity.